OpenStudy (nerdgirl):

Please help math problem two questions!!! Will fan and medal for quickest and correct answer! :) Is the relationship between the variables in the table a direct variation, an inverse variation, or neither? If it is a direct or inverse variation, write a function to model it. x 2 5 20 40 y 40 20 5 2 A. neither B. inverse variation; y = 20/x C. direct variation; y = 20x D. direct variation; y = x+15

OpenStudy (nerdgirl):

A drama club is planning a bus trip to New York City to see a Broadway play. The cost per person from the bus rental varies inversely as the number of people going on the trip. It will cost $30 per person if 44 people go on the trip. How much will it cost if 55 people go on the trip? A. $48.00 B. $12.50 C. $24.00 D. $33.00 @k_lynn @leah.desrosiers ANYONE!!!

OpenStudy (nerdgirl):

@Bigjuicydaddy @kyrabaaker @LeAnn01Xxx @H3LPN33DED @JulZak @Nnesha @donnie1999

OpenStudy (nerdgirl):

for which one?

OpenStudy (nerdgirl):

*sigh* Okay. I need help with the first one too though.

OpenStudy (nerdgirl):

@mathmath333 @mathmate @Red_hikari

OpenStudy (nerdgirl):

ANYONE!!!!!!!!!!!!!!!!!!!

OpenStudy (nerdgirl):

@misty1212

OpenStudy (nerdgirl):

@ganeshie8 @misty1212 @Nnesha

OpenStudy (anonymous):

hold on

OpenStudy (nerdgirl):

@misty1212 and ok

OpenStudy (misty1212):

HI!!

OpenStudy (nerdgirl):

hi can someone help me?

OpenStudy (misty1212):

which one are you doing?

OpenStudy (nerdgirl):

both. preferably the first one first then the second one @misty1212

OpenStudy (misty1212):

as x gets bigger, then y gets smaller right?

OpenStudy (nerdgirl):

yep

OpenStudy (misty1212):

lets check if it is \[y=20x\] is \[20=10\times 5\]?

OpenStudy (nerdgirl):

er....... sorry my brain is kind of fried normally i'd be able to do this quickly

OpenStudy (misty1212):

take your time,\(10\times 5\) is not that hard to do

OpenStudy (nerdgirl):

oh! right! sorry. hehe. dumb moment. 50

OpenStudy (misty1212):

ok so we can forget that one

OpenStudy (misty1212):

\[y=x+15\] is not direct variation so we can forget that one too

OpenStudy (nerdgirl):

so it's either a or b?

OpenStudy (misty1212):

and \[y=\frac{20}{x}\]means if you divide \(20\) by \(x\) you should get \(y\)

OpenStudy (misty1212):

if \(x=20\) then \(y=5\) from the table is \[5=\frac{20}{20}\]?

OpenStudy (nerdgirl):

20/20 is 1 so would it be b?

OpenStudy (misty1212):

?

OpenStudy (misty1212):

no not B since as you said \(\frac{20}{20}=1\) not \(5\)

OpenStudy (nerdgirl):

oh right! sorry. ugh....... I hate math. thanks again.

OpenStudy (misty1212):

yw

OpenStudy (nerdgirl):

can you help me with the second one too if i'm not too much of a pain?

OpenStudy (misty1212):

gotta do it quick \[y=\frac{k}{x}\] or \[xy=k\] you have \(y=30,x=44\) so \[k=30\times 44=1320\]

OpenStudy (misty1212):

put \(x=55\) and compute \[y=\frac{1320}{55}\] to get your answer good luckn !

OpenStudy (nerdgirl):

ty!